OpenStudy (anonymous):

Solve. Find the mean. Janice is selling cookies for her scout troop. One day she sold 10 boxes, the next day 15, the third day she sold 12, and on the fourth day she sold 13. How many does she need to sell on the 5th day to reach her goal of an average of 13 boxes a day?

OpenStudy (kirbykirby):

The average: (a+b+c+d+e)/5 Thus, (10+15+12+13+e)/5 = 13 (13 is the average) (50+e)/5 = 13 50+e = 13*5 50 +e = 65 e = 15 She needs to sell 15 boxes